OpenStudy (anonymous):

Classify: (Sqrroot 1/2 )(sqrroot 1/8 )

OpenStudy (anonymous):

is it \[\sqrt{1/2}) * \sqrt{1/8}\]

OpenStudy (anonymous):

Hey Nany ????????????

OpenStudy (anonymous):

Using the law \[\sqrt{m}*\sqrt{n} = \sqrt{mn}\] we get \[\sqrt{1/2}*\sqrt{1/8} = \sqrt{1/16}\]

OpenStudy (anonymous):

further using the law \[\sqrt{m/n} = \sqrt{m}/\sqrt{n}\] we get \[\sqrt{1}/\sqrt{16} = 1/4\]